Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

In a bowl of marbles, 8 are yellow, 6 are blue, and 4 are black. If Michelle picks 2 marbles out of the bowl at random a

Expert replies
by BTGmoderatorDC » Thu Sep 10, 2020 6:48 pm
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

In a bowl of marbles, 8 are yellow, 6 are blue, and 4 are black. If Michelle picks 2 marbles out of the bowl at random and at the same time, what is the probability that at least one of the marbles will be yellow?

(A) 5/17

(B) 12/17

(C) 25/81

(D) 56/81

(E) 4/9


OA B

Source: Veritas Prep
Join the discussion
Source: — Problem Solving |

BTGmoderatorDC wrote: ↑
Thu Sep 10, 2020 6:48 pm
In a bowl of marbles, 8 are yellow, 6 are blue, and 4 are black. If Michelle picks 2 marbles out of the bowl at random and at the same time, what is the probability that at least one of the marbles will be yellow?

(A) 5/17

(B) 12/17

(C) 25/81

(D) 56/81

(E) 4/9


OA B

Source: Veritas Prep
There is a total of 18 marbles.

It's more easier to start with the probability to pick none of the yellow marbles \(\Rightarrow \dfrac{10}{18} \cdot \dfrac{9}{17} = \dfrac{5}{17}\)

The probability to pick up at least one yellow marble (\(1-\) the probability to have none) will be
\(1 - \dfrac{5}{17} = \dfrac{12}{17} \Rightarrow\) B
Join the discussion

BTGmoderatorDC wrote: ↑
Thu Sep 10, 2020 6:48 pm
In a bowl of marbles, 8 are yellow, 6 are blue, and 4 are black. If Michelle picks 2 marbles out of the bowl at random and at the same time, what is the probability that at least one of the marbles will be yellow?

(A) 5/17

(B) 12/17

(C) 25/81

(D) 56/81

(E) 4/9


OA B

Solution:

We can look at this problem in terms of only 2 possible events. Either at least 1 yellow marble will be selected or no yellow marbles will be selected (but instead blue and black marbles will be selected). This means that

P(selecting at least 1 yellow marble) + P(selecting no yellow marbles) = 1

P(selecting at least 1 yellow marble) = 1 - P(selecting no yellow marbles)

Thus, if we can determine the probability that none of the 2 marbles selected are yellow, we’ll quickly be able to calculate the probability that at least one yellow marble is selected.

The bowl contains 8 marbles that are yellow, 6 that are blue, and 4 that are black, and we will select two of them. The probability that a non-yellow marble is selected first is 10/18 and the probability that a non-yellow marble is selected second is 9/17.

The probability that no yellow marbles are selected is: 10/18 x 9/17 = 5/1 x 1/17 = 5/17

Thus, the probability of selecting at least one yellow marble is: 1 - 5/17 = 12/17

Answer: B

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion